Calculate the pH of a solution formed by mixing 50.0mL of 0.40 M NH3 solution with 50.0mL of 0.40 M HCl solution. Is the resulting solution acidic, basic, or neutral? (Kb of NH3 = 1.8x10-5 ) (hint: determine if any reaction will occur between NH3 and HCl)

Can someone write this out neatly please
Answer #1

Given:

M(HCl) = 0.4 M

V(HCl) = 50 mL

M(NH3) = 0.4 M

V(NH3) = 50 mL

mol(HCl) = M(HCl) * V(HCl)

mol(HCl) = 0.4 M * 50 mL = 20 mmol

mol(NH3) = M(NH3) * V(NH3)

mol(NH3) = 0.4 M * 50 mL = 20 mmol

20 mmol of both will react to form NH4+ and H2O

NH3(aq) + HCl(aq) -> NH4+(aq) + Cl-(aq)

NH4+ here is acid.

So, the final solution will be acidic

Answer: acidic